OpenStudy (samigupta8):

In a ∆PQR is 3sinP+4cosQ=6 and 4sinQ+3cosP=1 then the angle R is equal to.... I m getting ans 5π/6 n π/6 Bt the ans is only one... How??

OpenStudy (samigupta8):

@parthkohli pls...help

OpenStudy (samigupta8):

@hartnn @ganeshie8 pls...help me

OpenStudy (faiqraees):

are you supposed to bring in a form Rsin/cos( x +theta)

OpenStudy (samigupta8):

No, we r required to find a solution of the two equations which gives us the value for P n Q ....n then find the value of R

OpenStudy (samigupta8):

Well, i get the final solution as Sin(P+Q)=1/2 this can be d case when P+Q=π/6 or 5π/6

OpenStudy (samigupta8):

So i m cuming up with value of R as π/6 or 5π/6 Bt the ans is only one out of d two....

OpenStudy (sparrow2):

for convenience: p=alpha ,q=beta. suppose alpha+beta=p/6